\begin{thebibliography}{10}

\bibitem{gunning2017explainable}
David Gunning.
\newblock Explainable artificial intelligence (xai).
\newblock {\em Defense advanced research projects agency (DARPA), nd Web},
  2(2):1, 2017.

\bibitem{Chromik2021}
Michael Chromik and Andreas Butz.
\newblock {Human-XAI Interaction: A Review and Design Principles for
  Explanation User Interfaces}.
\newblock {\em Lecture Notes in Computer Science (including subseries Lecture
  Notes in Artificial Intelligence and Lecture Notes in Bioinformatics)}, 12933
  LNCS:619--640, 2021.

\bibitem{liao2021human}
Q~Vera Liao and Kush~R Varshney.
\newblock Human-centered explainable ai (xai): From algorithms to user
  experiences.
\newblock {\em arXiv preprint arXiv:2110.10790}, 2021.

\bibitem{Abdul2018TrendsSystems}
Ashraf Abdul, Jo~Vermeulen, Danding Wang, Brian~Y. Lim, and Mohan Kankanhalli.
\newblock {Trends and Trajectories for Explainable, Accountable and
  Intelligible Systems}.
\newblock In {\em Proceedings of the 2018 CHI Conference on Human Factors in
  Computing Systems - CHI '18}, pages 1--18, 2018.

\bibitem{Doshi-Velez2017AInterpretability}
Finale Doshi-Velez and Been Kim.
\newblock {A roadmap for a rigorous science of interpretability}, 2017.

\bibitem{miller2019explanation}
Tim Miller.
\newblock Explanation in artificial intelligence: Insights from the social
  sciences.
\newblock {\em Artificial intelligence}, 267:1--38, 2019.

\bibitem{Zhu2018}
Jichen Zhu, Antonios Liapis, Sebastian Risi, Rafael Bidarra, and G.~Michael
  Youngblood.
\newblock {Explainable AI for Designers: A Human-Centered Perspective on
  Mixed-Initiative Co-Creation}.
\newblock {\em IEEE Conference on Computatonal Intelligence and Games, CIG},
  2018-Augus, 2018.

\bibitem{zhu2021player}
Jichen Zhu, Jennifer Villareale, Nithesh Javvaji, Sebastian Risi, Mathias
  L{\"o}we, Rush Weigelt, and Casper Harteveld.
\newblock Player-ai interaction: What neural network games reveal about ai as
  play.
\newblock In {\em Proceedings of the 2021 CHI Conference on Human Factors in
  Computing Systems}, pages 1--17, 2021.

\bibitem{geis2016cpux}
Thomas Geis, Knut Polkehn, Rolf Molich, and Oliver Kluge.
\newblock Cpux-ur curriculum.
\newblock {\em UXQB e. V}, 2016.

\bibitem{tintarev2007survey}
Nava Tintarev and Judith Masthoff.
\newblock A survey of explanations in recommender systems.
\newblock In {\em 2007 IEEE 23rd international conference on data engineering
  workshop}, pages 801--810. IEEE, 2007.

\bibitem{nunes2017systematic}
Ingrid Nunes and Dietmar Jannach.
\newblock A systematic review and taxonomy of explanations in decision support
  and recommender systems.
\newblock {\em User Modeling and User-Adapted Interaction}, 27(3):393--444,
  2017.

\bibitem{williams2021towards}
Oyindamola Williams.
\newblock Towards human-centred explainable ai: A systematic literature review.
\newblock 2021.

\bibitem{ehsan2021expanding}
Upol Ehsan, Q~Vera Liao, Michael Muller, Mark~O Riedl, and Justin~D Weisz.
\newblock Expanding explainability: Towards social transparency in ai systems.
\newblock In {\em Proceedings of the 2021 CHI Conference on Human Factors in
  Computing Systems}, pages 1--19, 2021.

\bibitem{liao2020questioning}
Q~Vera Liao, Daniel Gruen, and Sarah Miller.
\newblock Questioning the ai: informing design practices for explainable ai
  user experiences.
\newblock In {\em Proceedings of the 2020 CHI Conference on Human Factors in
  Computing Systems}, pages 1--15, 2020.

\bibitem{putnam2019exploring}
Vanessa Putnam and Cristina Conati.
\newblock Exploring the need for explainable artificial intelligence (xai) in
  intelligent tutoring systems (its).
\newblock In {\em IUI Workshops}, volume~19, 2019.

\bibitem{Tullio2007}
J~Tullio.
\newblock {\em How it works: A field study of non-technical users interacting
  with an intelligent system}.
\newblock Hale, 2007.

\bibitem{chen2021towards}
Zhongxia Chen, Xiting Wang, Xing Xie, Mehul Parsana, Akshay Soni, Xiang Ao, and
  Enhong Chen.
\newblock Towards explainable conversational recommendation.
\newblock In {\em Proceedings of the Twenty-Ninth International Conference on
  International Joint Conferences on Artificial Intelligence}, pages
  2994--3000, 2021.

\bibitem{dhurandhar2018explanations}
Amit Dhurandhar, Pin-Yu Chen, Ronny Luss, Chun-Chen Tu, Paishun Ting,
  Karthikeyan Shanmugam, and Payel Das.
\newblock Explanations based on the missing: Towards contrastive explanations
  with pertinent negatives.
\newblock {\em Advances in neural information processing systems}, 31, 2018.

\bibitem{Herlocker2000ExplainingRecommendations}
Jonathan~L. Herlocker, Joseph~A. Konstan, and John Riedl.
\newblock {Explaining collaborative filtering recommendations}.
\newblock {\em Proceedings of the 2000 ACM conference on Computer supported
  cooperative work - CSCW '00}, pages 241--250, 2000.

\bibitem{rader2018explanations}
Emilee Rader, Kelley Cotter, and Janghee Cho.
\newblock Explanations as mechanisms for supporting algorithmic transparency.
\newblock In {\em Proceedings of the 2018 CHI conference on human factors in
  computing systems}, pages 1--13, 2018.

\bibitem{ribeiro2016should}
Marco~Tulio Ribeiro, Sameer Singh, and Carlos Guestrin.
\newblock " why should i trust you?" explaining the predictions of any
  classifier.
\newblock In {\em Proceedings of the 22nd ACM SIGKDD international conference
  on knowledge discovery and data mining}, pages 1135--1144, 2016.

\bibitem{Stumpf2009}
Simone Stumpf, Vidya Rajaram, Lida Li, Weng~Keen Wong, Margaret Burnett, Thomas
  Dietterich, Erin Sullivan, and Jonathan Herlocker.
\newblock Interacting meaningfully with machine learning systems: Three
  experiments.
\newblock {\em International Journal of Human Computer Studies}, 67:639--662, 8
  2009.

\bibitem{Wang2019}
Danding Wang, Qian Yang, Ashraf Abdul, and Brian~Y. Lim.
\newblock {Designing theory-driven user-centric explainable AI}.
\newblock In {\em CHI '19 Proceedings of the 2019 annual conference on Human
  factors in computing systems}, 2019.

\bibitem{mohseni2021multidisciplinary}
Sina Mohseni, Niloofar Zarei, and Eric~D Ragan.
\newblock A multidisciplinary survey and framework for design and evaluation of
  explainable ai systems.
\newblock {\em ACM Transactions on Interactive Intelligent Systems (TiiS)},
  11(3-4):1--45, 2021.

\bibitem{cheng2019explaining}
Hao-Fei Cheng, Ruotong Wang, Zheng Zhang, Fiona O'Connell, Terrance Gray,
  F~Maxwell Harper, and Haiyi Zhu.
\newblock Explaining decision-making algorithms through ui: Strategies to help
  non-expert stakeholders.
\newblock In {\em Proceedings of the 2019 chi conference on human factors in
  computing systems}, pages 1--12, 2019.

\bibitem{narayanan2018humans}
Menaka Narayanan, Emily Chen, Jeffrey He, Been Kim, Sam Gershman, and Finale
  Doshi-Velez.
\newblock {\em arXiv preprint arXiv:1802.00682}, 2018.

\bibitem{selvaraju2017grad}
Ramprasaath~R Selvaraju, Michael Cogswell, Abhishek Das, Ramakrishna Vedantam,
  Devi Parikh, and Dhruv Batra.
\newblock Grad-cam: Visual explanations from deep networks via gradient-based
  localization.
\newblock In {\em Proceedings of the IEEE international conference on computer
  vision}, pages 618--626, 2017.

\bibitem{lipton2018mythos}
Zachary~C Lipton.
\newblock The mythos of model interpretability: In machine learning, the
  concept of interpretability is both important and slippery.
\newblock {\em Queue}, 16(3):31--57, 2018.

\bibitem{Bansal2018}
Gagan Bansal.
\newblock {Explanatory Dialogs: Towards Actionable, Interactive Explanations}.
\newblock In {\em AIES 2018 - Proceedings of the 2018 AAAI/ACM Conference on
  AI, Ethics, and Society}, pages 356--357, 2018.

\bibitem{norman1988psychology}
Donald~A Norman.
\newblock {\em The psychology of everyday things.}
\newblock Basic books, 1988.

\bibitem{Villareale2022}
Jennifer Villareale, Casper Harteveld, and Jichen Zhu.
\newblock {"I Want To See How Smart This AI Really Is”: Player Mental Model
  Development of an Adversarial AI Player}.
\newblock In {\em Proceedings of the 2022 ACM Conference of CHI PLAY}, 2022.

\bibitem{Kulesza2015PrinciplesLearning}
Todd Kulesza, Margaret Burnett, Weng-keen Wong, and Simone Stumpf.
\newblock {Principles of explanatory debugging to personalize interactive
  machine learning}.
\newblock In {\em Proceedings of the 20th international conference on
  intelligent user interfaces.}, pages 126--137, 2015.

\bibitem{Kulesza2013TooModels}
Todd Kulesza, Simone Stumpf, Margaret Burnett, Sherry Yang, Irwin Kwan, and
  Weng~Keen Wong.
\newblock {Too much, too little, or just right? Ways explanations impact end
  users' mental models}.
\newblock {\em Proceedings of IEEE Symposium on Visual Languages and
  Human-Centric Computing, VL/HCC}, pages 3--10, 2013.

\bibitem{Schaffer2015GettingAnalysis}
James Schaffer, Prasanna Giridhar, Debra Jones, Tobias H{\"{o}}llerer, Tarek
  Abdelzaher, and John O'donovan.
\newblock {Getting the message? A study of explanation interfaces for microblog
  data analysis}.
\newblock In {\em International Conference on Intelligent User Interfaces,
  Proceedings IUI}, volume 2015-Janua, pages 345--356, 2015.

\bibitem{hartson2003cognitive}
Rex Hartson.
\newblock Cognitive, physical, sensory, and functional affordances in
  interaction design.
\newblock {\em Behaviour \& information technology}, 22(5):315--338, 2003.

\end{thebibliography}

\documentclass{article}

   \usepackage[nonatbib]{neurips_2022}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         

\title{Towards Better User Requirements: How to Involve Human Participants in XAI Research}

\author{%
   Thu Nguyen \\
   IT University of Copenhagen \\
   Denmark \\
   \texttt{irng@itu.dk} \\
   \And
   Jichen Zhu \\
   IT University of Copenhagen \\
   Denmark \\
   \texttt{jichen.zhu@gmail.com} \\
}

\begin{document}

\maketitle

\begin{abstract}
Human-Center eXplainable AI (HCXAI) literature identifies the need to address user needs. This paper examines how existing XAI research involves human users in designing and developing XAI systems and identifies limitations in current practices, especially regarding how researchers identify user requirements. Finally, we propose several suggestions on how to derive better user requirements. 

\end{abstract}

\section{User Requirements for XAI Explanation}





The research field of eXplainable AI (XAI) has emerged to make AI more transparent and trustworthy to humans by opening the AI black-box and explaining its underlying operation\cite{gunning2017explainable}. While the field has made significant breakthroughs in {\em technical explainability}, it has limited success producing the {\em effective explanations} needed by users\cite{Chromik2021, liao2021human}. As a result, most explanations produced by XAI still lack usability, practical interpretability, and efficacy for real users~\cite{Abdul2018TrendsSystems, Doshi-Velez2017AInterpretability,miller2019explanation,Zhu2018,zhu2021player}. This viewpoint aligns with the one proposed by Liao and Kushney \cite{liao2021human}, where they argue that explanations should address stakeholders' needs. In our paper, we focus on understanding explanation needs from \textit{lay-end users}, who are direct users of the XAI system but have little domain and AI knowledge. This paper's {\em argument} is that a key step towards human-centered XAI and effective explanations is better-defined user requirements through deeper engagement with human users to gather that information. 

User requirements are insights from users about their needs, problems, and the context of use of an interactive system often derived from observing users performing tasks, interviewing users, and conducting focus groups \cite{geis2016cpux}. It is an essential part of the established User-Centered Design (UCD) process for interactive systems in general. Recently there has been growing acknowledgment that user requirements for explanation are necessary to build a usable XAI system \cite{liao2021human}. However, existing work typically attempted to define user requirements based on what the researchers identify as desired qualities of an XAI system \cite{tintarev2007survey, nunes2017systematic, williams2021towards}. While this approach helps to formalize the evaluation criteria, it risks bringing XAI researchers' bias as it is not coming from end users\cite{miller2019explanation}. Therefore, there is a need for methods to extract user requirements from end users directly. 



According to the established UCD process, user requirements should be derived before the XAI system's development to guide the development process towards an outcome that users will find valuable. However, in existing XAI research, only a small portion engages humans in deriving user requirements and other aspects of XAI development. Among them, many diverge from the established UCD design process regarding 1) when user requirements are collected, 2) how to collect user requirements for black-box models, and 3) which user groups are involved. The rest of this section will identify these issues, and the next section proposes approaches to address them.



With a few exceptions (e.g., \cite{ehsan2021expanding, liao2020questioning, putnam2019exploring, Tullio2007}), XAI development teams did not collect user requirements from human users at the early stage of XAI development yet. By contrast, a lot of existing research (e.g., \cite{chen2021towards, dhurandhar2018explanations, Herlocker2000ExplainingRecommendations, rader2018explanations, ribeiro2016should, Stumpf2009}) only engaged human users at the later stage to evaluate the XAI systems. This practice means that a lot of current XAI research is driven by researchers' intuition, and users are only asked at the end to validate the design decisions. 


Among these few projects that obtained user requirements, most seek to understand the user reasoning for Machine Learning models with high interpretability. For example, they focus on how users make sense of Naive Bayes \cite{Wang2019, Stumpf2009} and statistical modeling \cite{Tullio2007}, and use that information to derive user requirements. By contrast, there is a limited understanding of users' sense-making of black box models such as Deep Neural Networks. This is problematic because, without such knowledge, it is difficult to identify effective ways to communicate the information from XAI to users in ways that fit their cognitive needs.




Finally, user requirements gathering in current XAI research typically does not target specific lay-end user groups. As user requirements of explanation highly depend on the context of use \cite{geis2016cpux, liao2020questioning}, the user needs \cite{liao2021human, mohseni2021multidisciplinary}, and user profile \cite{geis2016cpux}. For example, AI developers need explanations to understand the inner working of the model to debug the algorithmic error \cite{liao2021human}. In contrast, non-expert end users will be overwhelmed by the same amount of technical details. For the same user group, whether they use AI in a high-stake or low-stake context will also affect the type of explanation they need. Also, their ability to understand explanations depends on their knowledge of the application domain and AI. Nonetheless, there are many XAI research studies (e.g., \cite{cheng2019explaining, narayanan2018humans, selvaraju2017grad, ribeiro2016should} rely on Amazon Mechanical Turk (AMT) workers, who do not represent a specific domain, needs, or context of use.  We argue that XAI user research should better consider the context of use, domain, and expertise of actual end users.  

\section{Proposed Approaches to Deriving Better HCXAI User Requirements}

 



We first suggest that human-centered XAI research should involve human users at the early stage of development. While there is growing recognition of the importance of the context of use and user needs \cite{liao2021human}, a common practice in the field is to derive user requirements from predefined desiderata such as \cite{lipton2018mythos, Bansal2018}. For example, Lipton~\cite{lipton2018mythos} proposed desiderata such as {\em trust, causality, transferability, informativeness, and fair and ethical decision making}. Bansal proposed {\em actionability} \cite{Bansal2018} as another desiderata for XAI. While these desiderata offer a general direction, they are often too generalized to target the specific requirements in a given context of use. For instance, what is considered trustworthy or informative can differ vastly from AI experts to novice end-users. Conducting user research, especially using qualitative methods (e.g., developing user group profiles, task models, and personas) at the early stage, can be a practical way to supplement the context-free desiderata.



Second, we argue that XAI user requirements should consider how to align users' mental models of the AI (i.e., how the user thinks the AI works) with the system's conceptual model (roughly speaking, how the AI works)\cite{norman1988psychology}. Existing HCI research has established that the usability of a traditional system improves when users' mental models align with the system's conceptual model. Due to the low human interpretability of black-box systems, however, it is unclear whether users, especially non-technical users, are able to construct mental models that completely match the system's conceptual model\cite{Villareale2022}. There are currently no agreements over whether XAI should include all details of system logic\cite{Kulesza2015PrinciplesLearning, Kulesza2013TooModels} or only selective important information\cite{Herlocker2000ExplainingRecommendations, Schaffer2015GettingAnalysis}). Researchers have identified cognitive science theories of how humans make sense of complex information as a potentially fruitful direction to design XAI systems\cite{Wang2019}. We believe that more users studies of how users construct mental models of AI  (e.g., \cite{Villareale2022}) and what cognitive-based design element (e.g., \textit{affordance} \cite{hartson2003cognitive}) can provide much-needed empirical evidence to bridge this knowledge gap. 

Lastly, XAI researchers should identify specific target user groups and develop XAI systems that address their requirements. Any real-world XAI system will be used by specific people (e.g., AI experts, domain experts, non-experts), in particular contexts (work, play, health), and with specific needs. This is especially true because the explanation is social interaction and needs to adapt to the explainees\cite{miller2019explanation}. To improve the usability and efficacy of XAI systems, researchers should identify the above key elements in user requirements. We believe that more research is needed in {\em explanation interfaces}\cite{Chromik2021,mohseni2021multidisciplinary} to study how to effectively {\em communicate} the technical information XAI algorithms generate to users. 

In conclusion, XAI research has yielded significant technical solutions to increase the interpretability and transparency of AI algorithms. To improve its real-world usefulness, we argue that more effort is needed to understand users of XAI systems, especially through better user requirements. 


\section{Acknowledgement}
This work is supported by the Danish Novo Nordisk Foundation Grant NNF20OC0066119.

\small{
\bibliographystyle{unsrt}
\bibliography{ms}}

\end{document}